# Towards Intelligent Control of MeV Energy Electrons and Protons from kHz Repetition Rate Ultra-Intense Laser Plasma Interactions

### v1

@@ II.2 Data Acquisition System
Our laser system is notable in that the control and data acquisition is a hybrid system that includes both EPICS-based controls and direct-connection controls. Fig. 2 shows a high level overview of this system. Originally the laser system operated purely with controls based on direct connection between the user interface computer and the device. This interface was limiting to our data acquisition needs, so we migrated the high data throughput instruments and important target and laser controls to EPICS. In this way, our laser system is an interesting example of modernizing a data acquisition and laser control system without completely re-designing it from the ground up. This hybrid approach provides two different ways to control and interface with the EPICS devices. This control is possible because LabView can be configured to send commands over EPICS and receive the data, or python scripts running independently of LabView can execute these commands. The former is useful for adjusting experimental conditions by hand while the latter is useful for allowing machine learning algorithms to direct the laser system to new parameters to explore as we will demonstrate in the next section.
@@ IV.2 Optimization
Figure 8 shows the log of the total electron CCD count for varying $\beta$ values with $KE_{\rm cutoff,goal}=0.7$ MeV. $KE_{\rm cutoff}=0.7$ MeV was chosen as the cutoff energy for this evaluation because each spectra has significant electron counts around 0.7 MeV. All models, except the 4-dimensional ridge regression, exhibit an increase in total electron counts as $\beta$ is varied from 1 to 0 as one would expect given the definition of Eq. LABEL:eq:function. Instead of showing detailed data for each model, we highlight some typical results from one of the models that performed well in Figure 8. Specifically, we show the electron energy spectra from the Gaussian process ($\alpha=2$) model. One can see that as $\beta$ decreases the total electron number counts increases significantly.
@@
Figure 8: Total electron CCD counts for different models while varying $\beta$ from 0 to 1 and keeping $KE_{\rm cutoff,goal}$ fixed at 0.7 MeV. The Ridge 4D model results were affected by target instability but the other models were not.
Figure 9: Electron energy spectra for the random forest regression with 100 estimators. Results shown are for $\beta=0$ to $\beta=1$ while keeping $KE_{\rm cutoff,goal}$ fixed at 0.7 MeV.
@@ V Discussion
This work was only possible due to significant improvements to the data acquisition system in our lab which included retrofitting our system with high throughput control using EPICS. This modification allowed for automated control of the parameters suggested via the optimization algorithm and data collection at a higher throughput rate. Single shot diagnostics of main pulse and pre-pulse energy were collected from over 5-million laser shots. The electron/ion spectrometer was improved to operate at 100 Hz but we collected at a rate of 10 Hz in order to increase the signal to noise. In total, over 50,000 unique electron and proton spectra were obtained for different experimental conditions, including different main pulse and pre-pulse energies, and target position. In this data, the time delay between the pre-pulse and the main pulse was fixed at 300 ps. In Appendix A, we present additional experimental data where the main pulse energy was fixed and the target held stationary near the peak focus, and where the pre-pulse energy and pre-pulse time delay were varied. To the best of our knowledge, this is one of the largest data sets for ultra-intense laser acceleration of electrons and protons ever collected.
@@
Another important aspect of our experiment was the “liquid-leaf" target system which is described well by George et al. (2019). A practical problem with these targets is that vibrations can lead to the breakup of the liquid sheet (e.g. Margarone et al. (2023)). Typically, studies with liquid leaf target systems involve relatively short timescales for data collection. For example, Treffert et al. (2022) produced deuteron and electron spectra for 60 consecutive laser shots in an experiment lasting around 2 minutes. By comparison, the data we present here typically involved running a liquid target system for about a half hour. We note that one of the models (Ridge 4D) was likely affected by the target stability, but we believe the other results to be unaffected by target stability concerns.
@@
During the optimization run, as highlighted in Figure 8, the number of electron counts does increase significantly with decreasing $\beta$ while keeping $KE_{\rm cutoff,goal}$ fixed at 0.7 MeV as one would expect given the definition of Eq. LABEL:eq:function. For context, Figure 9 shows detailed results for one of the models as $\beta$ was varied. Efforts to shift the maximum energy of the electron energy distribution had limited success (Figure 7). The placement of the electron spectrometer likely had a strong impact on that result. As shown in Figure 1, the electron spectrometer was placed in the target normal direction which raises the possibility that many electrons are being missed by the spectrometer. After these experiments were performed, we did briefly position the spectrometer in the forward laser direction, which increased the number of counts in the electron spectrometer but it did not increase the energies. We omit these results from the current paper for brevity. The electron spectrometer has detected higher electron energies in our lab previously. As described in Feister et al. (2017), when this same spectrometer was placed behind a hole in the OAP, electron energies were detected up to $\sim$2 MeV when the laser was near normal incidence on a liquid water column target. Clearly, there is interesting work to be done to explore different angles that electrons are ejected and which angles would be ideal for making a tunable MeV electron source. We left the spectrometer at target normal in this work because it is a dual electron/ion spectrometer and at these intensities the protons are being ejected from target normal.
@@
As mentioned earlier, the same experimental framework can be straightforwardly used to optimize and control proton acceleration. The ejected protons are likely more collimated than the electrons (c.f. single shot simulation results in Rahman et al. (2021)), so a challenge with performing analogous experiments with protons is the proton spectrometer reading would be very sensitive to the stability of the liquid leaf target. Efforts are currently being made to improve the stability of the target to allow for proton optimization experiments.
@@ Acknowledgements.
The data that support the findings of this study are available from the corresponding author upon reasonable request.
@@ Appendix A Time delay of the pre-pulse
This parameter scan lasted 40 minutes and it involved 2.4 million laser shots. Figure 11 shows the varied parameters along with the corresponding electron spectra during the duration of the run. Unlike the parameter scan described in Sec. IV, no significant proton signal was detected, so we do not include proton results in Figure. 11. As mentioned earlier, there is reason to believe that protons ejected from the laser interaction region are significantly more beamed than the electronsRahman et al. (2021). So even a small misalignment between the spectrometer and the perpendicular to the liquid sheet can lead to the proton beam missing the entry port to the spectrometer. We presume that this is the reason why negligible protons were detected.
@@
A notable result from this investigation, as one can see from Figure. 11, is that the pre-pulse time delay did affect the peak electron energies but at no point did the electron energies reach any higher than about 0.8 MeV, which is similar to what was found for other the other investigations in this paper. Figure 10 and 11 present measurements where the electron spectrometer is placed at target normal. As mentioned earlier in Sec. V, target normal may not be the best angle for the purpose of creating a tunable MeV electron source.

### v2

# Towards Intelligent Control of MeV Electrons and Protons from kHz Repetition Rate Ultra-Intense Laser Interactions

@@ II.2 Data Acquisition System
Our laser system is notable in that the control and data acquisition is a hybrid system that includes both EPICS-based controls and direct-connection controls. Fig. 2 shows a high level overview of this system. Originally the laser system operated purely with controls based on direct connection between the user interface computer and the device. This interface was limiting to our data acquisition needs. Our system was running with LabView in a synchronous manner and parameters could not be changed until all the data was saved. Firmware and software changes along with the introduction of EPICS allowed for the use of a common shot count across all data devices, which allows for asynchronous operation and data collection with the ability to combine the data for analysis with the common shot count. This improvement allows parameters to vary during an experiment while we know what changes correspond to different electron and ion spectra. We migrated important instruments and target and laser controls to EPICS. In this way, our laser system is an interesting example of modernizing a data acquisition and laser control system without completely re-designing it from the ground up. This hybrid approach provides two different ways to control and interface with the EPICS devices. This control is possible because LabView can be configured to send commands over EPICS and receive the data, or python scripts running independently of LabView can execute these commands. The former is useful for adjusting experimental conditions by hand while the latter is useful for allowing machine learning algorithms to direct the laser system to new parameters to explore as we will demonstrate in the next section.
@@ IV.2 Optimization
Figure 8 shows the log of the total electron (upper) and proton (lower) CCD count for varying electron $\beta$ values with $KE_{\rm cutoff,goal}=0.7$ MeV. $KE_{\rm cutoff}=0.7$ MeV was chosen as the cutoff energy for this evaluation because each spectra has significant electron counts around 0.7 MeV. All models, except the 4-dimensional ridge regression, exhibit an increase in total electron counts as $\beta$ is varied from 1 to 0 as one would expect given the definition of Eq. LABEL:eq:function. Likewise, the total proton counts increase with the electron counts as the electron $\beta$ is varied from 1 to 0. Instead of showing detailed data for each model, we highlight some typical results from one of the models that performed well in Figure 8. Specifically, we show the electron and proton energy spectra from the Gaussian process ($\alpha=2$) model. One can see that as $\beta$ decreases the total electron and proton number counts increase significantly.
@@
Figure 8: Total electron (upper) and proton (lower) CCD counts for different models while varying the electron $\beta$ from 0 to 1 and keeping $KE_{\rm cutoff,goal}$ fixed at 0.7 MeV. The Ridge 4D model results were affected by target instability but the other models were not.
Figure 9: Electron (upper) and proton (lower) energy spectra for the random forest regression with 100 estimators. Results shown are for $\beta=0$ to $\beta=1$ while keeping $KE_{\rm cutoff,goal}$ fixed at 0.7 MeV.
@@ V Discussion
This work was only possible due to significant improvements to the data acquisition system in our lab which included retrofitting our system with high throughput control using EPICS. This modification allowed for automated control of the parameters suggested via the optimization algorithm and data collection at a higher throughput rate. Single shot diagnostics of main pulse and pre-pulse energy were collected from over 5-million laser shots. The electron/ion spectrometer was improved to operate at 100 Hz but we collected at a rate of 10 Hz in order to increase the signal to noise. In total, over 50,000 unique electron and proton spectra were obtained for different experimental conditions, including different main pulse and pre-pulse energies, and target position. In this data, the time delay between the pre-pulse and the main pulse was fixed at 300 ps. In Appendix A, we present additional experimental data where the main pulse energy was fixed and the target held stationary near the peak focus, and where the pre-pulse energy and pre-pulse time delay were varied. This is an interesting parameter space to explore since varying the pre-pulse energy and timing will introduce different pre-plasma conditions to the target. Unfortunately, the data presented in Appendix A had significant target instability which hindered optimization efforts. The Appendix shows a proof of concept that time delay can be used in future optimization experiments with as many as 4 parameters. To the best of our knowledge, this is one of the largest data sets for ultra-intense laser acceleration of electrons and protons ever collected.
@@
Another important aspect of our experiment was the “liquid leaf" target system which is described well by George et al. (2019). A practical problem with these targets is that vibrations can lead to the breakup of the liquid sheet (e.g. Margarone et al. (2023)). Typically, studies with liquid leaf target systems involve relatively short timescales for data collection. For example, Treffert et al. (2022) produced deuteron and electron spectra for 60 consecutive laser shots in an experiment lasting around 2 minutes. By comparison, the data we present here typically involved running a liquid target system for about a half hour. We note that one of the models (Ridge 4D) was likely affected by the target stability, but we believe the other results to be unaffected by target stability concerns.
@@
During the optimization run, as highlighted in Figure 8, the number of electron and proton counts does increase significantly with decreasing $\beta$ while keeping $KE_{\rm cutoff,goal}$ fixed at 0.7 MeV as one would expect given the definition of Eq. LABEL:eq:function. For context, Figure 9 shows detailed results for one of the models as $\beta$ was varied. Efforts to shift the maximum energy of the electron energy distribution had limited success (Figure 7). The placement of the electron spectrometer likely had a strong impact on that result. As shown in Figure 1, the electron spectrometer was placed in the target normal rear direction which raises the possibility that many electrons are being missed by the spectrometer. After these experiments were performed, and at a time when the liquid jet system was producing thin sheets with heavy water (as also used by Knight et al. (2024)), we did briefly position the spectrometer in the forward laser direction. This spectrometer location showed a significant increased the number of electron counts and the electron kinetic energies increased to about 0.9 MeV with some electrons over 1 MeV. We did not perform optimizations with this configuration so we omit these results from the current paper for brevity. It bears mentioning that the electron spectrometer has detected even higher electron energies in our lab previously. As described in Feister et al. (2017), when this same spectrometer was placed behind a hole in the OAP, electron energies were detected up to $\sim$2 MeV when the laser was near normal incidence on a liquid water column target. Clearly, there is interesting work to be done to explore different angles that electrons are ejected and which angles would be ideal for making a tunable MeV electron source. We left the spectrometer at target normal in this work because it is a dual electron/ion spectrometer and at these intensities the protons are being ejected from target normal.
@@
As mentioned earlier, the same experimental framework can be straightforwardly used to optimize and control proton acceleration. The ejected protons are likely more collimated than the electrons (c.f. single shot simulation results in Rahman et al. (2021)), so a challenge with performing analogous experiments with protons is the proton spectrometer reading would be very sensitive to the stability of the liquid leaf target. In this experiment, the target suffered an instability that required adjustment 6 times. These instabilities are detrimental to the proton signal due to the signals narrow beam divergence. The electrons are accelerated in a much wider angle, which makes the electron signal more robust to the instabilities. Efforts are currently being made to improve the stability of the target to allow for proton optimization experiments.
@@ Acknowledgements.
The data that support the findings of this study are openly available zenodoTamminga (2025) at https://zenodo.org/records/15631826, reference number 10.5281/zenodo.15631826
@@ Appendix A Time delay of the pre-pulse
This parameter scan lasted 40 minutes and it involved 2.4 million laser shots. Figure 11 shows the varied parameters along with the corresponding electron spectra during the duration of the run. This parameter scan suffered from target instabilities which hindered the any optimization efforts using this dataset. Unlike the parameter scan described in Sec. IV, no significant proton signal was detected, so we do not include proton results in Figure. 11. As mentioned earlier, there is reason to believe that protons ejected from the laser interaction region are significantly more beamed than the electronsRahman et al. (2021). So even a small misalignment between the spectrometer and the perpendicular to the liquid sheet can lead to the proton beam missing the entry port to the spectrometer. We presume that this is the reason why negligible protons were detected.
@@
A notable result from this investigation (see Figure. 11) is that the pre-pulse time delay did affect the peak electron energies but at no point did the electron energies reach any higher than about 0.8 MeV, which is similar to what was found for other the other investigations in this paper. Figure 10 and 11 present measurements where the electron spectrometer is placed at target normal. As mentioned earlier in Sec. V, target normal may not be the best angle for the purpose of creating a tunable MeV electron source.